ON COMPUTATIONAL POWER OF QUANTUM BRANCHING PROGRAMS

FARID ABLAYEV AIDA GAINUTDINOVA MAREK KARPINSKI

Abstract. In this paper we study a model of a Quantum Branching Program (QBP) and investigate its computational power. We prove a general lower bound on the width of read-once QBPs, which we show to be almost tight on certain symmetric function.